At the end of the second day of the carnival some water was drained from the tank. the volume of water drained was 35.3 cubic feet. using your answer frm b detemine the height in feet of the water remaing in the tank after the water was drained. radius= 3 feet hieght is 5 feet but filled with 4 feet water answer from b is 75.3 cubic feet.
1
Expert's answer
2011-06-06T07:01:16-0400
Solution 1:
If V(from b)=75.3 cubic feet,
b=Vb/(πr2) = 75.3/(3.14*33)= 75.3/27.9=2.7 feet

Solution 2:
V(after drained) = V(tank) - V(drained) = πr2*h-35.3 = 3.14*32*4-35.3=76.3 cubic feet
b=(V(after drained)) / (πr2) = 76.3/ 27.9 = 2.7 feet